OpenStudy (anonymous):

Solve for X: Log 3 X = 2

OpenStudy (anonymous):

if it is log base 3 to 2 then x=3^2 i.e x =9

OpenStudy (anonymous):

you would have ti use what is called the definition of log: \[\Large \log_ab=x~~{\mathbb{is~~the~~same~~as~~writing}}~~b=a^x\]